A car radiator contains 10quarts of a 30% anitfreeze solution. How many quarts will have to be replaced with pure antifreeze if the resulting solution is to be 50% antifreeze
Answer by stanbon(75887)

A car radiator contains 10 quarts of a 30% anitfreeze solution. How many quarts will have to be replaced with pure antifreeze if the resulting solution is to be 50% antifreeze
------------------
Equation:
alcohol - alcohol + alcohol = alcohol
0.30*10 - 0.30x + 1.00x = 0.50*10
------------
Multiply thru by 100 to get:
30*10 - 30x + 100x = 50*10
70x = 20*10
x = (1/7)20
x = (20/7) quarts (amt. of 10% solution to replace by 100% antifreeze)
==================
Cheers,
Stan H.